OpenStudy (anonymous):

If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line? (2, 1) (1, 1) (0, 1)

OpenStudy (anonymous):

(2,1)

OpenStudy (anonymous):

no that is wrong sorry did not read it all

OpenStudy (anonymous):

slope is rise over run. I.e. chang in y/ change in x. so if you add one to the x componet you would have to add 2 to the y componet

OpenStudy (itsonlycdeee):

y= mx + b -1 = 0 ( 2) +b b = -1 ; therefore ...

OpenStudy (anonymous):

(0,-1) + (1,2) would give you (1,1) which is one of the points you listed

OpenStudy (anonymous):

thank u guys

OpenStudy (anonymous):

no problem. just remember (change in y)/(change in x) = slope and try to find some multiple of that fraction